DUE SOON, PLEASE HELP!!!!Sales for the final quarter of the prior year total 800 units. Expected sales in units) for the current year are: 720 (QuarteSales Budget Production Budget Direct Mtls Budget Direct Lbr Budget Factory OH Budget Selling Exp Budget Admin Exp Budget CosSales Budget Production Budget Direct Mtls Budget Direct Lbr Budget Factory OH Budget Selling Exp Budget Admin Exp Cost of In
Answer #1

Production Budget as follows:

Ham Inc.
Production Budget
For the year ended December 31, 2018
First Quarter Second Quarter Third Quarter Fourth Quarter Total
Next Quarter Budgeted Sales (Units) (A) 480 640 640 960
Ratio of Inventory to Future Sales (B) 50% 50% 50% 50%
Budgeted Ending Inventory (Units) (A × B)= C 240 320 320 480 480
Budgeted Sales (Units) (D) 720 480 640 640 2480
Required Units available production (E= C+D) 960 800 960 1120 2960
Budgeted beginning inventory (F = C, But the first quarter is given i.e. 360) 360 240 320 320 360
Units to be produced (G = E-F) 600 560 640 800 2600

____________________________________________________________________________________

Direct material budget as follows:

Ham Inc.
Direct materials Budget
For the year ended December 31, 2018
First Quarter Second Quarter Third Quarter Fourth Quarter Total
Units to be produced (G) 600 560 640 800
Material per unit (pounds) (H) 8 8 8 8
Materials needed for production (pounds) (I = G×H) 4800 4480 5120 6400
Add: Desired ending inventory (J = 50%(4480/5120/6400/3840 given) 2240 2560 3200 3840
Total material requirements (pounds) (K = I+J) 7040 7040 8320 10240
Less: Beginning Balance (L) 2400 2240 2560 3200
Materials to be purchased (pounds) (M= K-L) 4640 4800 5760 7040
Material Price-per pound (N) $14 $14 $14 $14
Total cost of direct materials purchases (O = M×N) $64,960 $67,200 $80,640 $98,560 $311,360